OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = . 5/6 What is the value of cot θ?

OpenStudy (merchandize):

sin θ = 5/6 sin^2 √ + cos^2 θ = 1 (5/6)^2 + cos^2 θ = 1 25/36 + cos^2 θ = 1 cos^2 θ = 11/36 cos θ = √11/6 cot θ = cos θ/sin θ = √11/6 ÷ 5/6 = √11/5

OpenStudy (anonymous):